Trigonometry question
If Cos47 + cos73 + cosx=0 .
Find the value of x.

Grade:8

3 Answers

Sooraj menon vk
7 years ago
using cos A+cos B=2*cos(A+B)/2*cos(A-B)/2
we have 2*cos60*cos 13.                                [cos(-B)=cosB]
or cos x+ cos 13=0,
cos x=cos(-13)
x=-13 degrees.
 
Seema
7 years ago
Answer for the   above  question is given   167
      
Suggestions  please.
Thanks
                  
  ‘’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’                               
            
abhisek
7 years ago
 
cos(47)+cos(73)
cos(30+17)+cos(90-17)
cos(30+17)+sin17                             according to     ASTC rule
cos(30).cos(17)-sin(30).sin(17)+sin(17)
cos(30).cos(17)-0.5.sin(17)+sin(17)
cos(30).cos(17)+0.5sin(17)
cos(30).cos(17)+sin(30).sin(17)
cos(30-17)=cos(13)
u can continue it from here